OpenStudy (pumpkin1102):

Variable a is 4 more than variable b. Variable a is also 1 less than b. Which of the following pairs of equations best models the relationship between a and b? a = b + 4 a = b − 1 a = 4b a = b + 1 b = a + 4 b = a − 1 b = 4a b = a + 1

OpenStudy (pumpkin1102):

I am 100% sure it isn't B

OpenStudy (texaschic101):

a is 4 more then b.... a = b + 4 u understand this one ?

OpenStudy (pumpkin1102):

I think so is it basically asking that when we solve for a we add 4 mor to a and that gives us the answer to B?

OpenStudy (shamim):

i think a=b+4 means a is greater than b right?

OpenStudy (pumpkin1102):

wouldn't it be C?

OpenStudy (shamim):

why u rsearching for result only

OpenStudy (pumpkin1102):

?

OpenStudy (texaschic101):

a = 1 less then b...how would u set this one up...any thoughts ?

OpenStudy (pumpkin1102):

1 - b = a

OpenStudy (texaschic101):

close...a = b - 1

OpenStudy (pumpkin1102):

o yeah I knew that

OpenStudy (pumpkin1102):

ok so now I think it's A or C but most likely A

OpenStudy (texaschic101):

so u have ur 2 equations...u know the answer ?

OpenStudy (pumpkin1102):

yay it's A

OpenStudy (texaschic101):

it is A

OpenStudy (pumpkin1102):

thx

OpenStudy (texaschic101):

yw